# Continual Learning of Large Language Models: A Comprehensive Survey

### v2

@@ 4.1.3. Distributional Shifts in CPT
Another pioneering work, LLPT (jin2022lifelong, ), establishes a comprehensive training and evaluation protocol for a series of content-level distributional shifts, referred to as “domain-incremental data streams”. They assess multiple CL methods and, similar to findings in (gogoulou2024continual, ), find consistent forward knowledge transfer, yet horizontal forgetting remains significant. Besides, contrary to the common understanding that experience replay (ER, (chaudhry2019tiny, )) is the most efficient approach to preventing forgetting, the authors find it ineffective in the case of CPT. They speculate that ER’s inefficiency may stem from overfitting issues (yuan2020revisiting, ; jin2022lifelong, ). Recyclable Tuning (qin2023recyclable, ) is the first study to consider both upstream LLM suppliers and downstream consumers at the same time. It shows that if the upstream supplier continually pre-trains LLMs, with or without replay, consumer-side efficiency can be boosted by recycling previously learned update components. Two CL techniques, initialization from outdated components and knowledge distillation, complement each other in improving adaptation efficiency in this context.
Other approaches involve training additional domain-specific experts for new content domains. DEMix (gururangan2022demix, ) addresses CPT by incrementally training and integrating new experts (DEMix layer) for new domains. To ensure reasonable inference performance during testing when no domain information is available, DEMix proposes a parameter-free probabilistic approach to dynamically estimate a weighted mixture of domains. Introducing a new domain variable $D_{t}$ alongside each word $x_{t}$, the authors estimate the next word probability $p(x_{t}|{\bm{x}}_{<t})$ by marginalizing over all experts [^2] In the marginalization step of the following equation in the original paper, it is $p(D_{t}=j|{\bm{x}}_{t})$ instead of $p(D_{t}=j|{\bm{x}}_{<t})$, while we believe this is a minor typo, and hence here we use the updated version.:
$p(x_{t}|{\bm{x}}_{<t})$ $=\sum_{j=1}^{n}p(x_{t}|{\bm{x}}_{<t},D_{t}=j)\cdot p(D_{t}=j|{\bm{x}}_{<t})=\sum_{j=1}^{n}p(x_{t}|{\bm{x}}_{<t},D_{t}=j)\cdot\left[\tfrac{p({\bm{x}}_{<t}|D_{t}=j)\cdot p(D_{t}=j)}{\sum_{j^{\prime}=1}^{n}p({\bm{x}}_{<t}|D_{t}=j^{\prime})\cdot p(D_{t}=j^{\prime})}\right],$
@@
where the conditional probability terms $p(\cdot|\cdot,D_{t})$ are calculated by using a specific domain expert. The DEMix framework’s modularization has been shown to facilitate efficient domain-adaptive pre-training, promote relevant knowledge during inference, and allow for removable components. Lifelong-MoE (chen2023lifelong, ), similar to DEMix (gururangan2022demix, ), incrementally trains domain experts for new domains. However, Lifelong-MoE differs from DEMix in utilizing a token-level gating function to activate multiple experts for intermediate embedding calculation. During training, previous experts’ parameters and gating functions remain frozen, and knowledge distillation loss is employed to regulate parameter updates, which thereby makes Lifelong-MoE robust against the issue of horizontal forgetting.
@@
Others. CPT as a technique to progressively attain novel knowledge, can be used to refine LLMs’ behavior. CEM (zhao2024large, ) collects examples where the model’s response is incorrect and continually trains the model on these, along with a supplemental dataset. RHO-1 (lin2024rho, ) proposes Selective Language Modeling (SLM), which employs a reference model to evaluate the perplexity of each token in the training corpus, and continually pre-trains the model on high-perplexity tokens. Similarly, IR-DRO (chen2024take, ) re-trains the model on re-weighted examples from the original pre-training dataset, focusing more on higher-loss sequences.
@@ 4.2. Domain-Adaptive Pre-training (DAP)
Summary of LLMs with DAP. We provide a summary of existing studies utilizing DAP for LLMs in Table 2. Each entry is characterized by three main features: (i) training process specifications, encompassing the vertical domain for which LLMs are trained, the training pipeline preceding release, and the LLM architecture employed; (ii) adopted continual learning techniques, including rehearsal, parameter regularization, and architecture expansion; and (iii) evaluation metrics for CL, such as backward transfer (forgetting) and forward transfer (adaptation to downstream data).
@@ 4.2.2. Different Domains of DAP
Legal Domain. Given the legal industry’s demand for managing ever-growing volumes of legal documents, there’s a burgeoning need to harness LLMs to aid legal professionals in navigating, interpreting, and generating high-quality legal materials (xiao2021lawformer, ; savelka2023explaining, ; yue2023disc, ). In Layer Llama (huang2023lawyer, ), the authors gathered publicly available legal texts from China Courts websites, totaling approximately 10 billion tokens as noted in a GitHub issue (lawyerllam-git, ). In SaulLM (colombo2024saullm7b, ), the authors collected the DAP corpus from various jurisdictions in different countries, resulting in a corpus of 30 billion tokens to cover diverse aspects of legal texts. When combined with previously available datasets (gao2020pile, ; koehn2005europarl, ), the total number of tokens used for legal-domain DAP reaches 94 billion. The substantial volume of DAP data, while offering valuable insights into specific domains, increases the risk of vertical forgetting of the general knowledge due to the large number of update steps involved. To mitigate this issue, SaulLM incorporates general data from Wikipedia, StackExchange, and GitHub into the DAP data, constituting about 2% of the final dataset (colombo2024saullm7b, ). Similarly, Lawyer Llama incorporates replaying general-domain data during DAP, but the replay rate is not disclosed (huang2023lawyer, ). (takahashi2024pretraining, ) also replays of non-latest business documents during DAP when building a Japanese business-specific LLM.
Medical Domain. The development of LLMs holds promise for revolutionary changes in the medical industry, offering potential improvements in efficiency and quality across medical communication, disease diagnosis, and decision-making for doctors (he2024foundation, ; li2023chatdoctor, ; singhal2023large, ; jeblick2022chatgpt, ; chen2023utility, ). Efforts have been made to develop medical specialists by either training an LLM from scratch (Han2023MedAlpaca, ; Singhal2023MedPaLM2, ; gu2021domain, ; luo2022biogpt, ) or fine-tuning publicly-available LLMs to meet specific medical needs (luo2023biomedgpt, ; wu2023pmc, ; Chen2023HuatuoGPTII, ; xiong2023doctorglm, ; bao2023discmedllm, ; zhang2023alpacare, ). Among these approaches, DAP techniques have been extensively utilized to preserve the communication and instruction-following abilities of a general LLM, preparing it for subsequent medical applications (luo2023biomedgpt, ; wu2023pmc, ; Chen2023HuatuoGPTII, ).
@@
BioMedGPT (luo2023biomedgpt, ) is a multi-modal biomedical language model that integrates representations of human language and the language of life (molecules, proteins, cells, genes, etc.). Prior to final multi-modal supervised fine-tuning, the authors initialize the model from Llama2-Chat (touvron2023llama2, ) and conduct DAP using extensive biomedical documents from S2ORC (lo2020s2orc, ), without considering any CL techniques or evaluations. In (guo2023continuous, ), DAP is performed using Chinese medical encyclopedias and online expert articles, with next-token prediction as the training objective. During DAP, the performance gradually deteriorates on general-domain datasets as the training step increases, but improves on the downstream medical examination tasks (hendryckstest2021, ; liu2023benchmarking, ; li2023cmmlu, ). PMC-LLama (wu2023pmc, ) gathers biomedical papers from S2ORC (lo2020s2orc, ) and medical textbooks for “knowledge injection training”. During this phase, a general language corpus from RedPajama-Data (together2023redpajama, ) is replayed at a 5% rate within a training batch. However, the paper does not analyze the effectiveness of this operation of mixing in general-domain data for DAP.
@@
To mitigate vertical forgetting, AF Adapter (yan2023af, ) proposes an adapter structure extending the width of Attention layers and FFNs for acquiring domain knowledge and only the adapters are tuned during DAP. Similarly, Hippocrates (acikgoz2024hippocrates, ) deploys LoRA during DAP to both have medical-specific knowledge injected and general ability preserved. Me-Llama (xie2024me, ) mixes in about 25% of the general-domain data for DAP on the clinical notes and biomedical articles, which achieves even positive backward transfer on MMLU (hendryckstest2021, ). HuatuoGPT-II (Chen2023HuatuoGPTII, ) proposes to fuse the DAP into the final SFT, making the two-stage development one unified process. The challenge of such process mainly comes from the data heterogeneity of DAP’s unlabeled corpus. The authors address this challenge by reformulating paragraphs of data into (instruction, output) format using existing large language models. They further employ a priority sampling strategy to avoid compromising downstream ability, a pitfall observed in the fixed-rate data mixing strategy (touvron2023llama2, ). This paper empirically demonstrates the superiority of unified one-stage SFT over two-stage training, questioning the reasonability of the current DAP. On medical-domain data, (rongali2021continual, ) finds that LMs constrained by CL techniques on source domains exhibit greater robustness to future domain shifts. Specifically, they identify that parameter regularization techniques like EWC (kirkpatrick2017overcoming, ), despite slightly higher cost, can facilitate positive forward and backward transfer.
@@
Financial Domain. Similar to the medical domain, LLMs hold immense potential for enhancing financial communication, decision-making processes, and risk assessment for both traders and ordinary individuals (shah2023zero, ; Yang2023InvestLM, ; Wang2023FinGPT, ; li2023cfgpt, ). Despite advancements, a gap persists between general-purpose LLMs and existing domain-specific smaller-scale LLMs (araci2019finbert, ; Wu2023BloombergGPT, ), underscoring the urgent need for more powerful financial-domain experts through the integration of LLMs. Notably, DAP techniques have emerged as crucial tools for tailoring LLMs to the intricacies of the financial domain while mitigating the negative effects of abrupt domain shifts from general to finance (Lu2023BBTFin, ; li2023cfgpt, ; xie2023efficient, ; Xue2023WeaverBird, ; Zhang2023xuanyuan, ).
@@
In (xie2023efficient, ), the authors aim to enhance the data efficiency of DAP. When the downstream tasks’ data distribution ${\mathcal{T}}$ are known, based on the generalization bound (ben2010theory, ; ganin2016domain, ; shi2024unified, ), the authors propose to sample the subset of DAP data whose distribution ${\mathcal{D}}$ is similar to the downstream task’s data, i.e., $d_{{\mathcal{H}}\Delta{\mathcal{H}}}({\mathcal{D}},{\mathcal{T}})$ is low. When the downstream data distribution is unknown, the authors suggest ensuring novelty and diversity in the sampled corpus for DAP. This approach significantly enhances DAP efficiency: it utilizes only 10% of the originally collected data yet outperforms models trained on the entire DAP dataset, underscoring the importance of data quality over quantity. WeaverBird (Xue2023WeaverBird, ) introduces an intelligent finance dialogue system, where the encoder is trained on Chinese and English financial documents, alongside expert-annotated financial query-response pairs, using LoRA (hu2022lora, ). Xuanyuan 2.0 (Zhang2023xuanyuan, ), akin to HuatuoGPT-II (Chen2023HuatuoGPTII, ), proposes the technique of hybrid-tuning, which fuses the stage of DAP and SFT into one, general-domain data and financial-domain data into one. Notably, the distribution of data in hybrid-tuning is non-conventional: financial DAP data comprises only a small portion of 13%. This prompts a pertinent question in line with the investigation on efficient DAP in (xie2023efficient, ): Is a large DAP dataset necessary for developing a domain-specific LLM?
Scientific Domain. Vertical scientific LLMs (Taylor2022Galactica, ; Yin2023FORGE, ; Xie2023DARWIN, ; Zhang2024SciGLM, ) span many subjects, including astronomy (Nguyen2023AstroLLaMA, ; Perkowski2024AstroLLaMAChat, ), mathematics (Azerbayev2023LLEMMA, ; Yu2023Outcome, ; luo2023wizardmath, ; yue2023mammoth, ; Gou2023ToRA, ), geology (Gao2023GLLAVA, ; roberts2023gpt4geo, ; lin2023geogalactica, ; wang2023nearrealtime, ), chemistry and physics (Rubungo2023LLM-Prop, ), biology (Yang2023PLLaMa, ; Zhao2023GIMLET, ; Cao2023InstructMol, ; Abdine2023Prot2Text, ; xTrimoPGLM2024Chen, ; Bi2023OCEANGPT, ; Zheng2023MarineGPT, ). However, among all the studies listed above, only a small fraction of them adopt the technique of DAP.
@@
OceanGPT (Bi2023OCEANGPT, ) is the first LLM tailored specifically for the ocean domain. It performs DAP on a raw corpus of ocean science literature, prioritizing recent research and historically significant works. K2 (deng2023learning, ) pioneers the development of a foundational language model tailored specifically for geoscience. It aggregates geoscience open access literature and Earth science-related Wikipedia pages for DAP. Following this, it undergoes multi-task instruction tuning utilizing LoRA (hu2022lora, ) on both a general instruction tuning dataset and the GeoSignal benchmark introduced within the K2 framework. AstroLlama (Nguyen2023AstroLLaMA, ) gathers abstracts solely from astronomy papers on arXiv and proceeds pre-training. It observes an improved perplexity on the domain of scholarly astronomy, without providing more quantitative evaluation. MarineGPT (Zheng2023MarineGPT, ) is a multi-modal LLM designed specifically for the marine domain. During DAP, MarineGPT incorporates 5 million marine image-text pairs to imbue domain knowledge. This involves training the parameters of a Q-Former (li2023blip2, ) between the frozen visual encoder (dosovitskiy2020image, ) and text decoder (touvron2023llama, ).
@@
Another branch of methods proactively integrate in the replay of the general-domain data to mitigate vertical forgetting. GeoGalactica (lin2023geogalactica, ) introduces a series of LLMs tailored for geoscience. In the DAP phase, besides the 52-billion-token geoscience corpus, Arxiv papers and Codedata are incorporated, with a mixing ratio of 8:1:1. The authors believe that the inclusion of the Codedata during the model’s pre-training can significantly boost the reasoning ability of the LLMs. Although GeoGalactica pinpoints challenges of DAP, including overfitting, catastrophic forgetting, maintaining the training stability, and convergence speed, it does not further provide empirical evidence supporting the inclusion of the Codedata, or deploying specific measures to address the challenges proposed above. Llemma (Azerbayev2023LLEMMA, ) focuses on mathematics, initialized from Code Llama (rozière2024code, ), and undergoes DAP on a blend of the 55 billion mathematical pre-training dataset and general domain data at the ratio of 19:1. In contrast, PLlama (Yang2023PLLaMa, ), designed for plant science, mixes domain-specific and general-domain data at the ratio of 9:1.
@@
Code Domain. The development of LLMs for automatic code filling, debugging, and generation holds significant practical importance (moradidakhel2023github, ; sun2024survey, ). These advancements cover various frameworks, including encoder-only (moradidakhel2023github, ), encoder-decoder (wang2021codet5, ; wang2023codet5plus, ; chai2023erniecode, ), and decoder-only (nijkamp2022codegen, ; nijkamp2023codegen2, ; zheng2023codegeex, ; chen2021evaluating, ; li2023starcoder, ; lozhkov2024starcoder, ; guo2024deepseekcoder, ). In the era of LLMs, there’s a growing trend towards decoder-only architectures (sun2024survey, ), leveraging models pre-trained on general natural language like Llama (touvron2023llama, ; touvron2023llama2, ). Consequently, there’s a shift in the training objective from utilizing code structures to simpler tasks like next token prediction and infilling.
@@
From the perspective of CL, the code domain presents unique advantages and challenges for DAP, compared to other domains discussed so far. On one hand, its hierarchical structure (general domain corpus $\rightarrow$ multi-language code $\rightarrow$ specific programming language) provides an ideal training pipeline for DAPs (rozière2024code, ), offering potential for more efficient training strategies. On the other hand, programming languages adhere to strict grammars, unlike the fuzzy and context-dependent natural language. Consequently, language models should ideally leverage these structures through tailored designs, and adopting the same training objectives as for natural languages may yield sub-optimal results. Therefore, many existing studies omit DAP (wang2021codet5, ; wang2023codet5plus, ; luo2023wizardcoder, ; muennighoff2024octopack, ; jiang2023selfevolve, ; wei2023magicoder, ; zhuo2024astraios, ; di2023codefuse, ; li2024instructcoder, ). In the following section, we will introduce existing code LLMs that employ DAP before the final downstream tasks, discussing both their common attributes and unique characteristics.
@@
Representing a series of notable works that focus solely on adaptation to target domains, CodeGen (nijkamp2022codegen, ) comprises a suite of LLMs designed for natural language (CodeGen-NL), multi-lingual programming languages (CodeGen-Multi), and mono-lingual programming languages (CodeGen-Mono). These models are trained sequentially, with each subsequent model initialized from the previous one trained on more general-domain data. Comment-Aug (song2024code, ) addresses the challenge of aligning programming languages with natural languages (PL-NL alignment) by performing DAP on the code augmented with generated additional comments. StarCoder (li2023starcoder, ) introduces two models: StarCoderBase and StarCoder. StarCoderBase is initially trained on a mixed dataset comprising various programming languages without significant reweighting on the data. Subsequently, StarCoderBase undergoes further fine-tuning on an additional 35 billion tokens of Python code, resulting in the development of StarCoder. DeepSeek-Coder-v1.5 (guo2024deepseekcoder, ) originates from DeepSeek-LLM (deepseekai2024deepseek, ) and undergoes pre-training on 2 trillion tokens, comprising 87% source code, 10% English code-related natural language, and 3% Chinese natural language corpus. Initialization from a general-domain LLM results in improved performance across various tasks, including natural language and mathematical reasoning, with minimal performance degradation on coding tasks, which underscores the efficacy of DAP.
@@
As the only work investigated so far that utilizes the general data replay to mitigate vertical forgetting in the code domain, Code Llama (rozière2024code, ) introduces a sophisticated training framework tailored for various coding tasks and model sizes. Initialized from Llama 2 weights, these models undergo DAP on a dataset composed of deduplicated public code, discussions about code, and a subset of natural language data. This mix of natural language data serves as a form of pseudo-replay to maintain the models’ proficiency in understanding natural language. Besides replay, architecture expansion has proven effective in acquiring robust coding abilities and preventing vertical forgetting simultaneously. IRCoder (paul2024ircoder, ) utilizes compiler intermediate representations to enhance the multilingual transferability of Code LLMs. By conducting DAP on code grounded in intermediate representations with LoRA (hu2021lora, ), IRCoder achieves superior multilingual programming instruction following, enhanced multilingual code understanding, and increased robustness to prompt perturbations. Llama Pro (wu2024llama, ) undergoes DAP on a combination of code and math data. It expands the original Llama2 architecture by dynamically adding multiple identity copies of the transformer blocks. These added blocks initially preserves the original functionality, and will be tuned for DAP. The proposed expansion method is shown to be more resilient against vertical forgetting compared to other parameter-efficient tuning methods like LoRA.

### v3

@@ 4.1.3. Distributional Shifts in CPT
Content Shift.(yildiz2024investigating, ) explores the large-scale CPT over 159 content domains, and shows that CPT on various domains can effectively improve models’ adaptation ability compared to DAP on single domain. Similarly, (gupta2023continual, ) continues the pre-training phase of Pythia (biderman2023pythia, ) with no complex CL techniques and discovers that learning rate re-warming consistently improves models trained from scratch. Built upon this simple observation, (ibrahim2024simple, ) further shows that proper combination of learning rate re-warming and re-decay, and replay of the previous data is sufficient to achieve a comparable performance to full re-training. LLPT (jin2022lifelong, ) establishes a comprehensive training and evaluation protocol for a series of content-level distributional shifts. They assess multiple CL methods and, similar to (gogoulou2024continual, ), find consistent forward knowledge transfer, yet horizontal forgetting remains significant. Besides, contrary to the common understanding that experience replay (chaudhry2019tiny, ) is the most efficient approach to preventing forgetting, the authors find it ineffective in the case of CPT, due to the potential overfitting issue. Recyclable Tuning (qin2023recyclable, ) shows that if the upstream supplier continually pre-trains LLMs, with or without replay, consumer-side efficiency can be boosted by recycling previously learned update components when proper CL techniques are applied.
@@
DEMix (gururangan2022demix, ) incrementally trains and integrates new experts (DEMix layer) for new domains during CPT. To ensure reasonable inference performance during testing when no domain information is available, it proposes a parameter-free probabilistic approach to dynamically estimate a weighted mixture of domains. DEMix’s modularization has been shown to facilitate efficient domain-adaptive pre-training, promote relevant knowledge during inference, and allow for removable components. Lifelong-MoE (chen2023lifelong, ), similar to DEMix (gururangan2022demix, ), incrementally trains domain experts for new domains. However, Lifelong-MoE differs from DEMix in utilizing a token-level gating function to activate multiple experts for intermediate embedding calculation. During training, previous experts’ parameters and gating functions remain frozen, and knowledge distillation loss is employed to regulate parameter updates, which thereby makes Lifelong-MoE robust against the issue of horizontal forgetting.
@@
Others. CPT as a technique to progressively attain novel knowledge, can be used to refine LLMs’ behavior. CEM (zhao2024large, ) collects examples where the model’s response is incorrect and continually trains the model on these examples, along with a supplemental dataset. RHO-1 (lin2024rho, ) proposes Selective Language Modeling (SLM), which employs a reference model to evaluate the perplexity of each token in the training corpus, and continually pre-trains the model on high-perplexity tokens. Similarly, IR-DRO (chen2024take, ) re-trains the model on re-weighted examples from the original pre-training dataset, focusing more on higher-loss sequences.
@@ 4.2. Domain-Adaptive Pre-training (DAP)
Summary of LLMs with DAP. We provide a summary of the existing 41 studies utilizing DAP for LLMs in Table 2. Each entry is characterized by three main features: (i) training process specifications, encompassing the vertical domain for which LLMs are trained, the training pipeline preceding release, and the LLM architecture employed; (ii) adopted continual learning techniques, including rehearsal, parameter regularization, and architecture expansion; and (iii) evaluation metrics for CL, such as backward transfer (forgetting) and forward transfer (adaptation to downstream data).
@@ 4.2.2. Different Domains of DAP
We include work aimed at establishing vertical LLMs across various domains, including legal, medical, financial, scientific, and code. Additionally, we cover other domains such as language and e-commerce.
@@
Legal Domain. In Layer Llama (huang2023lawyer, ), the authors gathered publicly available legal texts from China Courts websites, totaling approximately 10 billion tokens as noted in a GitHub issue. In SaulLM (colombo2024saullm7b, ), the authors collected the DAP corpus from various jurisdictions in different countries, resulting in a corpus of 30 billion tokens to cover diverse aspects of legal texts. When combined with previously available datasets, the total number of tokens used for legal-domain DAP reaches 94 billion. The substantial volume of DAP data, while offering valuable insights into specific domains, increases the risk of vertical forgetting of the general knowledge due to the large number of update steps involved. To mitigate this issue, SaulLM incorporates general data from Wikipedia, StackExchange, and GitHub into the DAP data, constituting about 2% of the final dataset (colombo2024saullm7b, ). Similarly, Lawyer Llama incorporates replaying general-domain data during DAP, but the replay rate is not disclosed (huang2023lawyer, ). (takahashi2024pretraining, ) also replays of non-latest business documents during DAP when building a Japanese business-specific LLM.
@@
Medical Domain. Efforts have been made to develop medical specialists by either training an LLM from scratch (gu2021domain, ; luo2022biogpt, ) or fine-tuning publicly-available LLMs to meet specific medical needs (luo2023biomedgpt, ; wu2023pmc, ; Chen2023HuatuoGPTII, ). Among these approaches, DAP techniques have been extensively utilized to preserve the communication and instruction-following abilities of a general LLM, preparing it for subsequent medical applications (luo2023biomedgpt, ; wu2023pmc, ; Chen2023HuatuoGPTII, ). BioMedGPT (luo2023biomedgpt, ) is a multi-modal biomedical language model that integrates representations of human language and the language of life (molecules, proteins, cells, genes, etc.). Prior to final multi-modal supervised fine-tuning, the authors initialize the model from Llama2-Chat (touvron2023llama2, ) and conduct DAP using extensive biomedical documents from S2ORC (lo2020s2orc, ), without considering any CL techniques or evaluations. In (guo2023continuous, ), DAP is performed using Chinese medical encyclopedias and online expert articles, with next-token prediction as the training objective. During DAP, the performance gradually deteriorates on general-domain datasets as the training step increases, but improves on the downstream medical examination tasks (hendryckstest2021, ). PMC-LLama (wu2023pmc, ) gathers biomedical papers from S2ORC (lo2020s2orc, ) and medical textbooks for “knowledge injection training.” During this phase, a general language corpus from RedPajama-Data (together2023redpajama, ) is replayed at a 5% rate within a training batch. However, the paper does not analyze the effectiveness of this operation of mixing in general-domain data for DAP.
@@
To mitigate vertical forgetting, AF Adapter (yan2023af, ) proposes an adapter structure extending the width of Attention layers and FFNs for acquiring domain knowledge and only the adapters are tuned during DAP. Similarly, Hippocrates (acikgoz2024hippocrates, ) deploys LoRA during DAP to both have medical-specific knowledge injected and general ability preserved. Me-Llama (xie2024me, ) mixes in about 25% of the general-domain data for DAP on the clinical notes and biomedical articles, which achieves even positive backward transfer on MMLU (hendryckstest2021, ). HuatuoGPT-II (Chen2023HuatuoGPTII, ) proposes to fuse the DAP into the final SFT, unifying the two stages into one single process. The challenge of such process mainly comes from the data heterogeneity of DAP’s unlabeled corpus. The authors address this challenge by reformulating paragraphs of data into (instruction, output) format using existing large language models. They further employ a priority sampling strategy to avoid compromising downstream ability, a pitfall observed in the fixed-rate data mixing strategy (touvron2023llama2, ). This paper empirically demonstrates the superiority of unified one-stage SFT over two-stage training, questioning the reasonability of the current DAP. On medical-domain data, (rongali2021continual, ) finds that LMs constrained by CL techniques on source domains exhibit greater robustness to future domain shifts. Specifically, they identify that parameter regularization techniques like EWC (kirkpatrick2017overcoming, ), despite slightly higher cost, can facilitate positive forward and backward transfer.
@@
Financial Domain. A gap persists between general-purpose LLMs and existing domain-specific smaller-scale LLMs (araci2019finbert, ; Wu2023BloombergGPT, ), underscoring the urgent need for more powerful financial-domain experts through the integration of LLMs. Notably, DAP techniques have emerged as crucial tools for tailoring LLMs to the intricacies of the financial domain while mitigating the negative effects of abrupt domain shifts from general to finance (Lu2023BBTFin, ; li2023cfgpt, ; xie2023efficient, ; Xue2023WeaverBird, ; Zhang2023xuanyuan, ).
@@
In (xie2023efficient, ), the authors aim to enhance the data efficiency of DAP. When the downstream tasks’ data distribution ${\mathcal{T}}$ are known, based on the generalization bound (ben2010theory, ; ganin2016domain, ; shi2024unified, ), the authors propose to sample the subset of DAP data whose distribution ${\mathcal{D}}$ is similar to the downstream task’s data, i.e., $d_{{\mathcal{H}}\Delta{\mathcal{H}}}({\mathcal{D}},{\mathcal{T}})$ is low. When the downstream data distribution is unknown, the authors suggest ensuring novelty and diversity in the sampled corpus for DAP. This approach significantly enhances DAP efficiency: it utilizes only 10% of the originally collected data yet outperforms models trained on the entire DAP dataset, underscoring the importance of data quality over quantity. WeaverBird (Xue2023WeaverBird, ) introduces an intelligent finance dialogue system, where the encoder is trained on Chinese and English financial documents, alongside expert-annotated financial query-response pairs, using LoRA (hu2022lora, ). Xuanyuan 2.0 (Zhang2023xuanyuan, ), akin to HuatuoGPT-II (Chen2023HuatuoGPTII, ), proposes the technique of hybrid-tuning, which fuses the stages of DAP and SFT into one, general-domain data and financial-domain data into one. Notably, the distribution of data in hybrid-tuning is unconventional: financial DAP data comprises only a small portion of 13%. This prompts a pertinent question in line with the investigation on efficient DAP in (xie2023efficient, ): Is a large DAP dataset necessary for developing a domain-specific LLM?
@@
Scientific Domain. Vertical scientific LLMs span many subjects (Zhang2024SciGLM, ; Nguyen2023AstroLLaMA, ; Azerbayev2023LLEMMA, ; luo2023wizardmath, ; lin2023geogalactica, ; Bi2023OCEANGPT, ; Zheng2023MarineGPT, ). However, among all the studies listed above, only a small fraction of them adopt the technique of DAP. OceanGPT (Bi2023OCEANGPT, ) is the first LLM tailored specifically for the ocean domain. It performs DAP on a raw corpus of ocean science literature, prioritizing recent research and historically significant works. K2 (deng2023learning, ) pioneers the development of a foundational language model tailored specifically for geoscience. It aggregates geoscience open access literature and Earth science-related Wikipedia pages for DAP. Following this, it undergoes multi-task instruction tuning utilizing LoRA (hu2022lora, ) on both a general instruction tuning dataset and the GeoSignal benchmark introduced within the K2 framework. AstroLlama (Nguyen2023AstroLLaMA, ) gathers abstracts solely from astronomy papers on arXiv and proceeds pre-training. It observes an improved perplexity on the domain of scholarly astronomy, without providing more quantitative evaluation. MarineGPT (Zheng2023MarineGPT, ) is a multi-modal LLM designed specifically for the marine domain. During DAP, MarineGPT incorporates 5 million marine image-text pairs to imbue domain knowledge. This involves training a Q-Former (li2023blip2, ) between the frozen visual and text decoder (dosovitskiy2020image, ; touvron2023llama, ).
@@
Another branch of methods proactively integrate in the replay of the general-domain data to mitigate vertical forgetting. GeoGalactica (lin2023geogalactica, ) introduces a series of LLMs tailored for geoscience. In the DAP phase, besides the 52-billion-token geoscience corpus, Arxiv papers and Codedata are incorporated, with a mixing ratio of 8:1:1. The authors believe that the inclusion of the Codedata during the model’s pre-training can significantly boost the reasoning ability of the LLMs. Although GeoGalactica pinpoints challenges of DAP, including overfitting, catastrophic forgetting, maintaining the training stability, and convergence speed, it does not further provide empirical evidence supporting the inclusion of the Codedata, or deploying specific measures to address the challenges proposed above. Llemma (Azerbayev2023LLEMMA, ) focuses on mathematics, initialized from Code Llama (rozière2024code, ), and undergoes DAP on a blend of the 55-billion-token mathematical pre-training dataset and general domain data at the ratio of 19:1. In contrast, PLlama (Yang2023PLLaMa, ), designed for plant science, mixes domain-specific and general-domain data at the ratio of 9:1.
@@
Code Domain. The development of LLMs for automatic code filling, debugging, and generation holds significant practical importance (moradidakhel2023github, ; sun2024survey, ). These advancements cover various frameworks, including encoder-only (moradidakhel2023github, ), encoder-decoder (wang2021codet5, ; wang2023codet5plus, ), and decoder-only (nijkamp2022codegen, ; lozhkov2024starcoder, ; guo2024deepseekcoder, ). There is a growing trend towards decoder-only architectures (sun2024survey, ), leveraging models pre-trained on general natural language like Llama (touvron2023llama, ; touvron2023llama2, ). Consequently, there is a shift in the training objective from utilizing code structures to simpler tasks like next token prediction and infilling.
@@
From the perspective of CL, the code domain presents unique advantages and challenges for DAP, compared to other domains. On one hand, its hierarchical structure (general domain corpus $\rightarrow$ multi-language code $\rightarrow$ specific programming language) provides an ideal training pipeline for DAPs (rozière2024code, ), offering potential for more efficient training strategies. On the other hand, programming languages adhere to strict grammars, unlike the fuzzy and context-dependent natural language. Consequently, language models should ideally leverage these structures through tailored designs, and adopting the same training objectives as for natural languages may yield sub-optimal results. Therefore, many existing studies omit DAP (wang2021codet5, ; wang2023codet5plus, ; luo2023wizardcoder, ). In the following section, we will introduce existing code LLMs that employ DAP before the final downstream tasks, discussing both their common attributes and unique characteristics.
@@
Representing a series of notable works that focus solely on adaptation to target domains, CodeGen (nijkamp2022codegen, ) comprises a suite of LLMs designed for natural language (CodeGen-NL), multi-lingual programming languages (CodeGen-Multi), and mono-lingual programming languages (CodeGen-Mono). These models are trained sequentially, with each subsequent model initialized from the previous one trained on more general-domain data. Comment-Aug (song2024code, ) addresses the challenge of aligning programming languages with natural languages (PL-NL alignment) by performing DAP on the code augmented with generated additional comments. StarCoder (li2023starcoder, ) introduces two models: StarCoderBase and StarCoder. StarCoderBase is initially trained on a mixed dataset comprising various programming languages without significant reweighting on the data. Subsequently, StarCoderBase undergoes further fine-tuning on additional 35 billion tokens of Python code, resulting in the development of StarCoder. DeepSeek-Coder-v1.5 (guo2024deepseekcoder, ) originates from DeepSeek-LLM (deepseekai2024deepseek, ) and undergoes pre-training on 2 trillion tokens, comprising 87% source code, 10% English code-related natural language, and 3% Chinese natural language corpus. Initialization from a general-domain LLM results in improved performance across various tasks, including natural language and mathematical reasoning, with minimal performance degradation on coding tasks, which underscores the efficacy of DAP.
@@
As the only work that utilizes the general data replay to mitigate vertical forgetting in the code domain, Code Llama (rozière2024code, ) introduces a sophisticated training framework tailored for various coding tasks and model sizes. Initialized from Llama 2 weights, these models undergo DAP on a dataset composed of deduplicated public code, discussions about code, and a subset of natural language data. This mix of natural language data serves as a form of pseudo-replay to maintain the models’ proficiency in understanding natural language. Besides replay, architecture expansion has proven effective in acquiring robust coding abilities and preventing vertical forgetting simultaneously. IRCoder (paul2024ircoder, ) utilizes compiler intermediate representations to enhance the multilingual transferability of Code LLMs. By conducting DAP on code grounded in intermediate representations with LoRA (hu2021lora, ), IRCoder achieves superior multilingual programming instruction following, enhanced multilingual code understanding, and increased robustness to prompt perturbations. Llama Pro (wu2024llama, ) undergoes DAP on a combination of code and math data. It expands the original Llama2 architecture by dynamically adding multiple identity copies of the transformer blocks. These added blocks initially preserves the original functionality, and will be tuned for DAP. The proposed expansion method is shown to be more resilient against vertical forgetting compared to other parameter-efficient tuning methods like LoRA.